If 3x-y=12, then y/3 =


Answer by unlockmath(1688)

Hello,
For this one:
3x-y=12
Subtract 3x from both sides to get:
-y=-3x+12
Now divide by -3 to get:
y/3=x-4
Make sense?
RJ
www.math-unlock.com